# RanDumb: A Simple Approach that Questions the Efficacy of Continual Representation Learning

### v1

@@ 2 RanDumb: Mechanism & Intuitions
EWC [24] (PNAS ’17) 0 10.1$\pm$ 0.0 10.6$\pm$ 0.4 1.0$\pm$ 0.0 1.0$\pm$ 0.0
SI [60] (ICML ’17) 0 12.7$\pm$ 1.0 10.1$\pm$ 0.1 1.1$\pm$0.0 1.0$\pm$0.1
LwF [26] (TPAMI ’17) 0 11.8 $\pm$ 0.6 10.1$\pm$ 0.1 0.9$\pm$0.0 1.0$\pm$ 0.0
LT [61] (Arxiv ’18) 0 10.9$\pm$ 0.9 10.0$\pm$ 0.2 1.1$\pm$ 0.1 1.0$\pm$ 0.0
Gen-NCM [22] (NeurIPS-W ’22) 0 82.0$\pm$ 0.0 27.7$\pm$ 0.0 10.0$\pm$ 0.0 7.5$\pm$ 0.0
Gen-SLDA [20] (CVPR-W ’20) 0 88.0$\pm$ 0.0 41.4$\pm$ 0.0 18.8$\pm$ 0.0 12.9$\pm$ 0.0
VAE-GC [50] (CVPR-W ’21) 0 84.0$\pm$ 0.5 42.7$\pm$ 1.3 19.7$\pm$ 0.1 12.1$\pm$ 0.1
PEC [59] (ICLR ’24) 0 92.3$\pm$ 0.1 58.9$\pm$ 0.1 26.5$\pm$ 0.1 14.9$\pm$ 0.1
RanDumb (Ours) 0 98.3 (+5.9) 55.6 (-3.3) 28.6 (+2.1) 17.7 (+2.8)
3 Experiments
We compare RanDumb with algorithms across online continual learning benchmarks with an emphasis on exemplar-free and low-exemplar storage regime complementing GDumb [39] in regimes where it has poor performance.
Benchmarks. We illustrate the benchmarks used in our experiments along with key differences to training run for RanDumb in Table 3. We aim for a truly comprehensive coverage and show results on four different benchmarks (A, B, D, E) which reflect the latest trends in online continual learning (’22-’24) across exemplar-free, contrastive-training [^2] Benchmark B is split into two distinct sections: (i) methods that do not rely on contrastive learning and heavy augmentation, named B1, and (ii) approaches that incorporate contrastive learning and extra augmentations, named as Scenario B2., meta-continual learning and network-expansion based approaches respectively and a rehearsal-free offline continual learning benchmark C. These benchmarks are ordered by increasingly relaxed constraints, moving further away from the training scenario of RanDumb. These benchmarks train models from scratch. Hence, RanDumb utilizes a random Fourier transform for embedding.
We further test on exemplar-free scenarios in offline continual learning using Benchmark F [54] with the challenging one-class per task constraint borrowed from [59]. This benchmark allows using pretrained models along with unrestricted training time and access to all class samples at each timestep. However, RanDumb is restricted to learning from a single pass seeing only one sample at a time. RanDumb only learns a linear classifier over a given pretrained model in Benchmark F. We use LAMDA-PILOT codebase [^3] https://github.com/sun-hailong/LAMDA-PILOT for all methods, except RanPAC and SLDA which use their codebase. All methods use the original hyperparameters. We only change initial classes to 2 and number of classes per task to 1 in the scripts and test using both ImageNet21K and ImageNet1K ViT-B/16 models. For detailed assumptions, task orderings, implementation details and other specific information of these benchmarks and approaches, readers are encouraged to consult the LAMDA-PILOT codebase.
Implementation Details (RanDumb). We evaluate RanDumb using five datasets commonly used for online continual learning: MNIST, CIFAR10, CIFAR100, TinyImageNet200, and miniImageNet100. For the latter two datasets, we downscale all images to 32x32. We augment each datapoint with flipped version, hence two images are seen by the classifier at each timestep (except for MNIST and Benchmark F). We normalize all images and flatten them into vectors, obtaining 784-dim input vectors for MNIST and 3072-dim input vectors for all the other datasets. For Benchmark F, we compare RanDumb on seven datasets used in LAMDA-PILOT, replacing ObjectNet with Stanford Cars as ObjectNet license prohibits training models on it. We use the 768-dimensional features from the same pretrained ViT-B models used in this benchmark. We measure accuracy on the test set of all past seen classes after completing the full one-pass over the dataset. We take the average accuracy after the last task on all past tasks [59, 18, 54]. In Benchmark A and F, since we have one class per task, the average accuracy across past tasks is the same regardless of the task ordering. In Benchmarks A-E, all datasets have the same number of samples, hence similarly the average accuracy across past tasks is the same regardless of the task ordering. We used the Scikit-Learn implementation of Random Fourier Features [42] with 25K embedding size, $\gamma=1.0$. We use progressively increasing ridge regression parameter ($\lambda$) with dataset complexity, using $\lambda=10^{-6}$ for MNIST, $\lambda=10^{-5}$ for CIFAR10/100 and $\lambda=10^{-4}$ for TinyImageNet200/miniImageNet100. All experiments were conducted on a CPU server with a 48-core Intel Xeon Platinum 8268 CPU and 392GB of RAM. Our code is available at https://github.com/drimpossible/RanDumb.
Table 5: Benchmark B.1 (Ref: Table adopted from OnPro [55], OCM[18]) We perform comprehensive comparisons of RanDumb with popular online continual learning approaches in less ideal many-classes per task setting referred as ‘Dataset (num_tasks/num_classes_per_task)’. Furthermore, results are categorized by memory buffer sizes given in ‘M’ subcolumn for that dataset. Comparing across rows, we observe that RanDumb outperforms the best among the compared approaches without heavy-augmentations by 3-20% margins despite being exemplar free. Only in one case, it is second best, but after GDumb. Comparing the two columns of CIFAR100, we observe a large drop in performance of OCL methods on benchmarks with longer timesteps indicating analyzing longer timesteps is important in continual learning.
Method MNIST (5/2) CIFAR10 (5/2) CIFAR100 (10/10) CIFAR100 (50/2) TinyImageNet (100/2)
$M=0.1k$ $M=0.1k$ $M=0.2k$ $M=0.5k$ $M=1k$ $M=1k$ $M=1k$ $M=2k$
AGEM [12] (ICLR ’19) 56.9±5.2 17.7±0.3 22.7±1.8 5.8±0.2 5.9±0.1 1.8±0.2 0.8±0.1 0.9±0.1
GSS [3] (NeurIPS ’19) 70.4±1.5 18.4±0.2 26.9±1.2 8.1±0.2 11.1±0.2 4.3±0.2 1.1±0.1 3.3±0.5
ER [13] (ICML-W ’19) 78.7±0.4 19.4±0.6 29.7±1.0 8.7±0.3 15.7±0.3 8.3±0.3 1.2±0.1 5.6±0.5
ASER [46] (AAAI ’21) 61.6±2.1 20.0±1.0 27.8±1.0 11.0±0.3 16.4±0.3 9.6±1.3 2.2±0.1 5.3±0.3
MIR [2] (NeurIPS ’19) 79.0±0.5 20.7±0.7 37.3±0.3 9.7±0.3 15.7±0.2 12.7±0.3 1.4±0.1 6.1±0.5
ER-AML [10] (ICLR ’22) 76.5±0.1 - 40.5±0.7 - 16.1±0.4 - - 5.4±0.2
iCaRL [44] (CVPR ’17) - 31.0±1.2 33.9±0.9 12.8±0.4 16.5±0.4 - 5.0±0.3 6.6±0.4
DER++ [9] (NeurIPS ’20) 74.4±1.1 31.5±2.9 44.2±1.1 16.0±0.6 21.4±0.9 9.3±0.3 3.7±0.4 5.1±0.8
GDumb [39] (ECCV ’20) 81.2±0.5 23.3±1.3 35.9±1.1 8.2±0.2 18.1±0.3 18.1±0.3 4.6±0.3 12.6±0.1
CoPE [15] (CVPR ’21) - 33.5±3.2 37.3±2.2 11.6±0.4 14.6±1.3 - 2.1±0.3 2.3±0.4
DVC [18] (CVPR ’22) - 35.2±1.7 41.6±2.7 15.4±0.3 20.3±1.0 - 4.9±0.6 7.5±0.5
Co²L [11] (ICCV ’21) 83.1±0.1 - 42.1±1.2 - 17.1±0.4 - - 10.1±0.2
R-RT [6] (CVPR ’21) 89.1±0.3 - 45.2±0.4 - 15.4±0.3 - - 6.6±0.3
CCIL [35] (CVPR ’21) 86.4±0.1 - 50.5±0.2 - 18.5±0.3 - - 5.6±0.9
IL2A [65] (NeurIPS ’21) 90.2±0.1 - 54.7±0.5 - 18.2±1.2 - - 5.5±0.7
BiC [56] (CVPR ’19) 90.4±0.1 - 48.2±0.7 - 21.2±0.3 - - 10.2±0.9
SSIL [1] (ICCV ’21) 88.2±0.1 - 49.5±0.2 - 26.0±0.1 - - 9.6±0.7
Rehearsal-Free
PASS [65] (NeurIPS ’21) - 33.7±2.2 33.7±2.2 7.5±0.7 7.5±0.7 - 0.5±0.1 0.5±0.1
@@ 3 Experiments
Table 6: (Left) Benchmark B.2 (Ref: Table from OnPro [55]) We now compare with state-of-the-art contrastive representation learning based online continual learning approaches which additionally use sophisticated augmentations. These augmentations and additional loss function can improve any of the OCL methods described in Setup B.1 by large margins [18]. We observe that RanDumb often outperforms these sophisticated methods despite all of these factors on small-exemplar settings. (Right) Benchmark C (Ref: Table 2 from [48]). We compare the performance of RanDumb with latest efforts on improving rehearsal-free methods. We outperform them by 4% margins.
@@
RanDumb 28.6 (+4.0)
Table 7: (Left) Benchmark D (Ref: Table 2 from VR-MCL [4]) We compare RanDumb with meta-continual learning approaches operating in a high memory setting, allowing buffer sizes up to 1K exemplars. Despite catering representations specifically to combat forgetting, all methods with the exception of VR-MCL are outperformed by RanDumb on TinyImageNet. RanDumb also surpasses all prior work by a substantial 9.1% on CIFAR100. Allowing generous replay buffers shifts scenarios to a high exemplar regime where GDumb performs the best on CIFAR10. Yet RanDumb competes favorably even under these conditions not optimized for its approach. (Middle) Benchmark E (Ref: Table 1 from SEDEM [58]) We compare RanDumb with recent network expansion based online continual learning approaches. Despite allowing access to much larger memory buffers, RanDumb matches the performance of best method SEDEM on MNIST, while exceeding it by 0.3% on CIFAR10 and 3.8% on CIFAR100. (Right) Architectures (Ref: Table 1 from Mirzadeh et al. [34]) RanDumb surpasses continual representation learning across a wide range of architectures, achieving close to 94% of the joint performance.
@@
A-GEM [12] (ICLR ’19) 18.4 $\pm$ 0.2 6.0 $\pm$ 0.2 4.0 $\pm$ 0.2
IS [60] (ICML ’17) 17.4 $\pm$ 0.2 5.2 $\pm$ 0.2 3.3 $\pm$ 0.3
MER [45] (ICLR ’19) 36.9 $\pm$ 2.4
La-MAML [19] (NeurIPS ’20) 33.4 $\pm$ 1.2 11.8 $\pm$ 0.6 6.74 $\pm$ 0.4
GDumb [39] (ECCV ’20) 61.2 $\pm$ 1.0 18.1 $\pm$ 0.3 4.6 $\pm$ 0.3
ER [13] (ICML-W ’19) 43.8 $\pm$ 4.8 16.1 $\pm$ 0.9 11.1 $\pm$ 0.4
DER [9] (NeurIPS ’20) 29.9 $\pm$ 2.9 6.1 $\pm$ 0.1 4.1 $\pm$ 0.1
DER++ [9] (NeurIPS ’20) 52.3 $\pm$ 1.9 11.8 $\pm$ 0.7 8.3 $\pm$ 0.3
CLSER [5] (ICLR ’22) 52.8 $\pm$ 1.7 17.9 $\pm$ 0.7 11.1 $\pm$ 0.2
OCM [18] (ICML ’22) 53.4 $\pm$ 1.0 14.4 $\pm$ 0.8 4.5 $\pm$ 0.5
ER-OBC [14] (ICLR ’23) 54.8 $\pm$ 2.2 17.2 $\pm$ 0.9 11.5 $\pm$ 0.2
VR-MCL [4] (ICLR ’24) 56.5 $\pm$ 1.8 19.5 $\pm$ 0.7 13.3 $\pm$ 0.4
@@
MIR [2] (NeurIPS ’19) 93.2 $\pm$ 0.4 42.8 $\pm$ 2.2 20.0 $\pm$ 0.6
GEM [12] (ICLR ’19) 93.2 $\pm$ 0.4 24.1 $\pm$ 2.5 11.1 $\pm$ 2.4
iCARL [44] (CVPR ’17) 83.9 $\pm$ 0.2 37.3 $\pm$ 2.7 10.8 $\pm$ 0.4
G-MED [23] (NeurIPS ’21) 82.2 $\pm$ 2.9 47.5 $\pm$ 3.2 19.6 $\pm$ 1.5
GSS [3] (NeurIPS ’19) 92.5 $\pm$ 0.9 38.5 $\pm$ 1.4 13.1 $\pm$ 0.9
CoPE [15] (CVPR ’21) 93.9 $\pm$ 0.2 48.9 $\pm$ 1.3 21.6 $\pm$ 0.7
CURL [43] (NeurIPS ’19) 92.6 $\pm$ 0.7 - -
CNDPM [25] (ICLR ’20) 95.4 $\pm$ 0.2 48.8 $\pm$ 0.3 22.5 $\pm$ 1.3
Dynamic-OCM [57] (ECCV ’22) 94.0 $\pm$ 0.2 49.2 $\pm$ 1.5 21.8 $\pm$ 0.7
SEDEM [58] (ICCV ’23) 98.3 $\pm$ 0.2 55.3 $\pm$ 1.3 24.8 $\pm$ 1.2
@@
RandDumb (Ours) 74.8 (-2.0)
3.1 Results
We extensively evaluate RanDumb and detail our findings on each benchmark in this section.
Benchmark A. Benchmark A assesses continual learning models in the challenging setup of one class per timestep, closely mirroring our training assumptions. We present our results in Table 4. Comparing across rows, and see that RanDumb improves over prior state-of-the-art across all datasets with 2-6% margins. The only exception is PEC on CIFAR10, where RanDumb underperforms by 3.3%. Nonetheless, it still outperforms the second-best model, GDumb with a 500 memory size, by 4.9%.
Benchmark B1. We present our results comparing with non-contrastive methods in Table 5. We notice that scenario allows two classes per task and relaxes the memory constraints for online continual learning methods, allowing for higher accuracies compared to Benchmark A. Despite that, RanDumb outperforms latest OCL algorithms on MNIST, CIFAR10 and CIFAR100—often by margins exceeding $>$10%. The lone exception is GDumb achieving a higher performance with 2K memory samples on TinyImageNet, indicating that this already is in the high-memory regime. Overall, RanDumb is the best algorithm across the benchmark.
Benchmark B2. We additionally compare our performance with the latest online CL approaches using contrastive losses with sophisticated data augmentations. As shown in in Table 6 (Left), these advancements provide large performance improvements over methods from Benchmark B.1. To compensate, we compare on lower exemplar budgets. The best approach, OnPro [55], which outpeforms RanDumb on CIFAR10 by 2.2% and TinyImageNet by 0.3%, but falls significantly short on CIFAR100 by 5.9%. Overall, RanDumb achieves strong results compared to highly performant representation learning using state-of-the-art contrastive learning approaches customized to continual learning, despite storing no exemplars.
Benchmark C. We now compare against offline rehearsal-free continual learning approaches in Table 6 (Right) on CIFAR100 dataset. Despite online training, RanDumb outperforms PredKD by over 4% margins.
Benchmark D. We now compare performance of RanDumb against meta-continual learning methods, which require large exemplars with buffer sizes of 1K in Table 7 (left). RanDumb achieves strong performance under these conditions, exceeding all prior work by a large margin of 9.1% on CIFAR100 and outperforms all but VR-MCL approach on the TinyImageNet dataset. GDumb performs the best on CIFAR10, indicating this is already in a large-exemplar regime uniquely unsuited for RanDumb.
Benchmark E. We compare RanDumb against recent network expansion-based online continual learning methods in Table 7 (middle). These approaches grow model capacity to mitigate forgetting while dealing with shifts in the data distribution, and are allow larger memory buffers. RanDumb matches the performance of the state-of-the-art method SEDEM [58] on MNIST, while exceeding it by 0.3% on CIFAR10 and 3.8% on CIFAR100. The results indicate current benchmarks remain too restrictive for effective continual representation learning.
Benchmark F. We compare performance of approaches which do not further train the deep network like RanDumb (ours) [^4] Recall that RanDumb in Benchmark F is quite similar to SLDA [20]. and NCM [22] against popular prompt-tuning approaches in Table 8. and discover that prompt-tuning approaches completely collapse under large timesteps and approaches which do not finetune their pretrained model achieve strong performance, even under challenging one class per timestep constraint. Note that there are variants of RanDumb (e.g. RanPAC - RP - PETL [30]) that achieve higher accuracies due to minor modifications which are designed for this benchmark.
Conclusion. Overall, despite RanDumb being exemplar-free and modeling worst-case ordering, it outperforms nearly all online continual learning methods across various tasks when exemplar storage is limited. We specifically benchmark on lower exemplar sizes to complement settings in which GDumb does not perform well.
Table 8: Benchmark F We compare RanDumb with other prompt-tuning based continual learning approaches using ViT-B/16 ImageNet-21K/1K pretrained models using 2 init classes and 1 class per task setting. Most prompt-tuning based methods collapse and simple baselines which do not pretrain the architecture like NCM [22] or RanDumb achieve state-of-the-art performance. ’-’ indicates that despite reasonable efforts, we could not run the method. ${}^{+}$here representation learning and random projection both collapse, reporting last stable accuracy.
@@ 3.1 Results
L2P [54] (CVPR ’22) 2.4 0.3 0.8 1.4 1.3
DualPrompt [53] (ECCV ’22) 2.3 0.3 0.8 0.9 4.2
CODA-Prompt [47] (CVPR ’23) 2.6 0.3 0.8 1.9 6.3
Adam-Adapt [64] (Arxiv ’23) 76.7 49.3 62.0 85.2 83.6
Adam-SSF [64] (Arxiv ’23) 76.0 47.3 64.2 85.6 84.2
Adam-VPT [64] (Arxiv ’23) 79.3 35.8 61.2 83.8 86.9
Adam-FT [64] (Arxiv ’23) 72.6 49.3 61.0 85.2 83.8
Memo [63] (ICLR ’23) 69.8 - - 81.4 -
iCARL [44] (CVPR ’17) 72.4 - 35.2 72.4 -
Foster [52] (ECCV ’22) 52.2 - 76.8 86.6 -
NCM [22] (NeurIPS-W ’22) 76.2 49.4 61.2 85.2 83.6
SLCA [62] (ICCV ’23) 86.3 - 52.8 84.7 -
RanPAC [30] (NeurIPS ’23) 88.2 ${}^{+}$ 72.8 ${}^{+}$ 93.0
RanDumb (Ours) 84.5 47.8 66.9 88.0 93.4
@@
Adam-Adapt [64] (Arxiv ’23) 82.4 48.8 55.4 86.7 84.4
Adam-SSF [64] (Arxiv ’23) 82.7 46.0 59.7 86.2 84.9
Adam-VPT [64] (Arxiv ’23) 70.8 34.8 53.9 84.0 81.1
Adam-FT [64] (Arxiv ’23) 65.7 48.5 56.1 86.5 84.4
Foster [52] (ECCV ’22) 87.3 - 5.1 86.9 -
iCARL [44] (CVPR ’17) 71.6 - 35.1 71.6 -
NCM [22] (NeurIPS-W ’22) 81.3 48.9 54.6 86.7 84.4
SLCA [62] (ICCV ’23) 86.8 - 54.2 82.1 -
RanPAC [30] (NeurIPS ’23) 89.6 ${}^{+}$ 67.3 ${}^{+}$ ${}^{+}$
RanDumb (Ours) 86.8 42.2 64.9 88.5 92.3
Table 9: RanDumb (Analysis). We thoroughly analyse the key components of RanDumb, and observe: (1) contributions of decorrelation and embedding, finding them interdependent and integral to performance; (2) consistent gains from augmentation across datasets; (3) marginal returns and saturation with higher embedding sizes, enabling computational tradeoffs; (4) increase in the optimal regularisation parameter for more complex datasets; (5) comparisons with alternate embeddings (or lack thereof) favoring RanDumb’s approach.
Method MNIST CIFAR10 CIFAR100 T-ImNet m-ImNet
(10/1) (10/1) (10/1) (200/1) (100/1)
Ablating Components of RanDumb
RanDumb 98.3 55.6 28.6 11.1 17.7
-Embed 88.0 (-10.3) 41.6 (-14.0) 19.0 (-9.6) 8.0 (-3.1) 12.9 (-4.8)
-Both 82.1 (-16.2) 28.5 (-27.1) 10.4 (-18.2) 4.1 (-7.0) 7.28 (-10.4)
Effect of Adding Flip Augmentation
With - 55.6 28.6 11.1 17.7
Without 98.3 52.5 (-3.1) 26.9 (-1.7) 10.7 (-0.4) 16.6 (-1.1)
Variation with Ridge Parameter $\lambda$
$\lambda=10^{-6}$ 98.3 53.9 27.8 10.3 15.8
$\lambda=10^{-5}$ 55.6 28.6 11.1 15.9
$\lambda=10^{-4}$ 96.6 52.6 26.1 11.6 17.7
Variation Across Embedding Projections
No-Embed 88.0 41.6 19.0 8.0 12.9
RP+ReLU 95.2 48.8 23.1 9.7 15.7
3.2 RanDumb: Analysis
We extensively analyse RanDumb for the traditional OCL benchmarks in this subsection. First, we ablate the two components (random embedder and decorrelation) in RanDumb to ensure both components are necessary. Then we study effect of varying the embedding dimensionality, shrinkage and removing flipping augmentation and compare with alternate random embeddings proposed in literature [30] and across architectures [34].

### v2

@@ 3 Experiments
PASS [81] - 33.7±2.2 33.7±2.2 7.5±0.7 7.5±0.7 - 0.5±0.1 0.5±0.1
@@
Implementation Details (RanDumb). We evaluate RanDumb using five datasets: MNIST, CIFAR10, CIFAR100, TinyImageNet200, and miniImageNet100. For the latter two, we downscale all images to 32x32. We augment each datapoint with flipped version, hence two images are seen by the classifier at each timestep (except for MNIST and Benchmark F). We normalize all images and flatten them into vectors, obtaining 784-dim input vectors for MNIST and 3072-dim input vectors for all the other. For Benchmark F, we compare RanDumb on seven datasets used in LAMDA-PILOT, replacing ObjectNet with Stanford Cars as ObjectNet license prohibits training models. We use the 768-dimensional features from the same pretrained ViT-B models used in this benchmark. We measure accuracy on the test set of all past seen classes after completing the full one-pass. We take the average accuracy after the last task on all past tasks [75, 25, 67]. In Benchmark A and F, since we have one class per task, the average accuracy across past tasks is the same regardless of the task ordering. In Benchmarks A-E, all datasets have the same number of samples, hence similarly the average accuracy across past tasks is the same regardless of the task ordering. We used the Scikit-Learn implementation of Random Fourier Features [52] with 25K embedding size, $\gamma=1.0$. We use progressively increasing ridge regression parameter ($\lambda$) with dataset complexity, $\lambda=10^{-6}$ for MNIST, $\lambda=10^{-5}$ for CIFAR10/100 and $\lambda=10^{-4}$ for TinyImageNet200/miniImageNet100.
Table 5: (Left) Benchmark B.2 (Ref: Table from OnPro [68]) We compare RanDumb with contrastive representation learning based approaches which additionally use sophisticated augmentations. We observe that RanDumb often outperforms these sophisticated methods despite all of these factors on small-exemplar settings. (Right) Benchmark C (Ref: Table 2 from [60]). We compare RanDumb with latest rehearsal-free methods. RanDumb outperforms them by 4% margin.
@@
RanDumb (Ours) 28.6 (+4.0)
Table 6: (Left) Benchmark D (Ref: Table 2 from VR-MCL [71]) We compare RanDumb with meta-continual learning approaches operating in a high memory setting, allowing buffer sizes up to 1K exemplars. RanDumb outperforms all methods except VR-MCL on TinyImageNet. RanDumb also surpasses all prior work by a substantial 9.1% on CIFAR100. Allowing generous replay buffers shifts scenarios to a high exemplar regime where GDumb performs the best on CIFAR10. Yet RanDumb competes favorably even under these conditions. (Right) Benchmark E (Ref: Table 1 from SEDEM [74]) We compare RanDumb with network expansion based approaches. Despite allowing access to much larger memory buffers, RanDumb matches the performance of best method SEDEM on MNIST, while exceeding it by 0.3% on CIFAR10 and 3.8% on CIFAR100.
@@
LWF [37] 18.8 $\pm$ 0.1 5.6 $\pm$ 0.4 4.0 $\pm$ 0.3
A-GEM [12] 18.4 $\pm$ 0.2 6.0 $\pm$ 0.2 4.0 $\pm$ 0.2
IS [76] 17.4 $\pm$ 0.2 5.2 $\pm$ 0.2 3.3 $\pm$ 0.3
MER [55] 36.9 $\pm$ 2.4
La-MAML [26] 33.4 $\pm$ 1.2 11.8 $\pm$ 0.6 6.74 $\pm$ 0.4
GDumb [49] 61.2 $\pm$ 1.0 18.1 $\pm$ 0.3 4.6 $\pm$ 0.3
ER [13] 43.8 $\pm$ 4.8 16.1 $\pm$ 0.9 11.1 $\pm$ 0.4
DER [9] 29.9 $\pm$ 2.9 6.1 $\pm$ 0.1 4.1 $\pm$ 0.1
DER++ [9] 52.3 $\pm$ 1.9 11.8 $\pm$ 0.7 8.3 $\pm$ 0.3
CLSER [5] 52.8 $\pm$ 1.7 17.9 $\pm$ 0.7 11.1 $\pm$ 0.2
OCM [25] 53.4 $\pm$ 1.0 14.4 $\pm$ 0.8 4.5 $\pm$ 0.5
ER-OBC [18] 54.8 $\pm$ 2.2 17.2 $\pm$ 0.9 11.5 $\pm$ 0.2
VR-MCL [71] 56.5 $\pm$ 1.8 19.5 $\pm$ 0.7 13.3 $\pm$ 0.4
@@
MIR [3] 93.2 $\pm$ 0.4 42.8 $\pm$ 2.2 20.0 $\pm$ 0.6
GEM [12] 93.2 $\pm$ 0.4 24.1 $\pm$ 2.5 11.1 $\pm$ 2.4
iCARL [54] 83.9 $\pm$ 0.2 37.3 $\pm$ 2.7 10.8 $\pm$ 0.4
G-MED [32] 82.2 $\pm$ 2.9 47.5 $\pm$ 3.2 19.6 $\pm$ 1.5
GSS [4] 92.5 $\pm$ 0.9 38.5 $\pm$ 1.4 13.1 $\pm$ 0.9
CoPE [19] 93.9 $\pm$ 0.2 48.9 $\pm$ 1.3 21.6 $\pm$ 0.7
CURL [53] 92.6 $\pm$ 0.7 - -
CNDPM [36] 95.4 $\pm$ 0.2 48.8 $\pm$ 0.3 22.5 $\pm$ 1.3
Dynamic-OCM [73] 94.0 $\pm$ 0.2 49.2 $\pm$ 1.5 21.8 $\pm$ 0.7
SEDEM [74] 98.3 $\pm$ 0.2 55.3 $\pm$ 1.3 24.8 $\pm$ 1.2
@@
3.1 Results
Benchmark A. We assess continual learning models in the challenging setup of one class per timestep, closely mirroring our training assumptions, and present our results in Table 3. Comparing across rows, and see that RanDumb improves over prior state-of-the-art across all datasets with 2-6% margins. The only exception is PEC on CIFAR10, where RanDumb underperforms by 3.3%. Nonetheless, it outperforms the second-best model, GDumb with a 500 memory size, by 4.9%.
Benchmark B1. We present our results comparing with non-contrastive methods in Table 4. We notice that scenario allows two classes per task and relaxes the memory constraints for online continual learning methods, allowing for higher accuracies compared to Benchmark A. Despite that, RanDumb outperforms latest OCL algorithms on MNIST, CIFAR10 and CIFAR100—often by margins exceeding 10%. The lone exception is GDumb achieving a higher performance with 2K memory samples on TinyImageNet, indicating this already is in the high-memory regime.
Benchmark B2. We additionally compare our performance with the latest OCL approaches using contrastive losses with sophisticated data augmentations. As shown in in Table 5 (Left), these advancements provide large performance improvements over methods from Benchmark B.1. To compensate, we compare on lower exemplar budgets. The best approach, OnPro [68], outperforms RanDumb on CIFAR10 by 2.2% and TinyImageNet by 0.3%, but falls significantly short on CIFAR100 by 5.9%. Overall, RanDumb achieves strong results compared to representation learning using state-of-the-art contrastive learning approaches customized to continual learning, despite storing no exemplars.
Benchmark C. We compare against offline rehearsal-free continual learning approaches in Table 5 (Right) on CIFAR100. Despite online training, RanDumb outperforms PredKD by over 4% margins.
Benchmark D. We compare performance of RanDumb against meta-continual learning methods, which require large exemplars with buffer sizes of 1K in Table 6 (left). RanDumb achieves strong performance under these conditions, exceeding all prior work by a large margin of 9.1% on CIFAR100 and outperforms all but VR-MCL approach on the TinyImageNet dataset. GDumb performs the best on CIFAR10, indicating this is already in a large-exemplar regime uniquely unsuited for RanDumb.
Benchmark E. We compare RanDumb against network expansion-based online continual learning methods in Table 6 (right). These approaches grow model capacity to mitigate forgetting while dealing with shifts in the data distribution, and are allowed larger memory buffers. RanDumb matches the performance of the state-of-the-art method SEDEM [74] on MNIST, while exceeding it by 0.3% on CIFAR10 and 3.8% on CIFAR100.
Table 7: (Left) Analysis of RanDumb: We study contributions of decorrelation, random embedding, and data augmentation. We further vary the embedding sizes and regularisation parameter. Finally, we compare with alternate embeddings. (Right) Architectures (Ref: Table 1 from Mirzadeh et al. [45]) RanDumb surpasses continual representation learning across a wide range of architectures, achieving close to 94% of the joint performance.
Method MNIST CIFAR10 CIFAR100 T-ImNet m-ImNet
(10/1) (10/1) (10/1) (200/1) (100/1)
Ablating Components of RanDumb
RanDumb 98.3 55.6 28.6 11.1 17.7
-Embed 88.0 (-10.3) 41.6 (-14.0) 19.0 (-9.6) 8.0 (-3.1) 12.9 (-4.8)
-Both 82.1 (-16.2) 28.5 (-27.1) 10.4 (-18.2) 4.1 (-7.0) 7.28 (-10.4)
Effect of Adding Flip Augmentation
With - 55.6 28.6 11.1 17.7
Without 98.3 52.5 (-3.1) 26.9 (-1.7) 10.7 (-0.4) 16.6 (-1.1)
Variation with Ridge Parameter $\lambda$
$\lambda=10^{-6}$ 98.3 53.9 27.8 10.3 15.8
$\lambda=10^{-5}$ 55.6 28.6 11.1 15.9
$\lambda=10^{-4}$ 96.6 52.6 26.1 11.6 17.7
Variation Across Embedding Projections
No-Embed 88.0 41.6 19.0 8.0 12.9
RP+ReLU (RanPAC) 95.2 48.8 23.1 9.7 15.7
@@ 3.1 Results
RanDumb (Ours) 74.8 (-2.0)
Table 8: Benchmark F We compare RanDumb with prompt-tuning approaches using ViT-B/16 ImageNet-21K/1K pretrained models using 2 init classes and 1 class per task setting. Most prompt-tuning based methods collapse, RanDumb achieves either state-of-the-art or second-best performance. RanPAC-imp is an improved version of the RanPAC mitigating the instability issues.
@@
L2P [67] 2.4 0.3 0.8 1.4 1.3
DualPrompt [66] 2.3 0.3 0.8 0.9 4.2
CODA-Prompt [59] 2.6 0.3 0.8 1.9 6.3
Adam-Adapt [80]) 76.7 49.3 62.0 85.2 83.6
Adam-SSF [80] 76.0 47.3 64.2 85.6 84.2
Adam-VPT [80] 79.3 35.8 61.2 83.8 86.9
Adam-FT [80] 72.6 49.3 61.0 85.2 83.8
Memo [79] 69.8 - - 81.4 -
iCARL [54] 72.4 - 35.2 72.4 -
Foster [65] 52.2 - 76.8 86.6 -
NCM [31] 78.3 44.3 62.5 84.8 88.2
SLCA [78] 86.3 - 52.8 84.7 -
RanPAC [41] 88.2 39.0 72.8 77.7 93.0
RanPAC-imp [41] 87.8 43.5 72.6 89.6 93.0
RanDumb (Ours) 84.5 49.5 66.9 88.0 93.6
@@
Adam-Adapt [80] 82.4 48.8 55.4 86.7 84.4
Adam-SSF [80] 82.7 46.0 59.7 86.2 84.9
Adam-VPT [80] 70.8 34.8 53.9 84.0 81.1
Adam-FT [80] 65.7 48.5 56.1 86.5 84.4
Foster [65] 87.3 - 5.1 86.9 -
iCARL [54] 71.6 - 35.1 71.6 -
NCM [31] 83.5 41.4 54.8 86.5 88.5
SLCA [78] 86.8 - 54.2 82.1 -
RanPAC [41] 89.6 26.8 67.3 87.2 88.2
RanPAC-imp [41] 89.4 33.8 69.4 89.6 91.9
RanDumb (Ours) 86.8 42.2 64.9 88.5 92.4
Benchmark F. We compare performance of approaches which do not further train the deep network like RanDumb against popular continual finetuning and prompt-tuning approaches in Table 8. We discover that prompt-tuning approaches completely collapse under large timesteps and approaches which do not finetune their pretrained model achieve strong performance, even under challenging one class per timestep constraint. Note that RanPAC [41] adds a RP+ReLU and finetunes in a first-session adaptation fashion over RanDumb, yet fails to achieve higher accuracies.
Overall, despite RanDumb being exemplar-free, it outperforms nearly all online continual learning methods across various tasks when exemplar storage is limited. We specifically benchmark on lower exemplar sizes to complement settings in which GDumb does not perform well.
3.2 Analysis of RanDumb